# Spectral selectors and contact orderability

### v1

@@ 2.2. The $C^{1}$-topology of Legendrian and contact isotopy classes
Let $\Lambda_{*}$ be a closed Legendrian submanifold of $(M,\xi)$. Let us briefly describe the $C^{1}$-topology of the space $\mathcal{L}(\Lambda_{*})$ of Legendrian submanifolds isotopic to $\Lambda_{*}$ and the space $\mathcal{G}(M,\xi)$ of compactly supported contactomorphisms isotopic to the identity. The $C^{1}$-topology on $\mathcal{G}$ is understood as the topology induced by the Whitney $C^{1}$-topology on diffeomorphisms as it is described in [42, Chapter 2] (both weak and strong topologies coincide for compactly supported diffeomorphisms). One way to define the $C^{1}$-topology on $\mathcal{L}(\Lambda_{*})$ is to consider $\mathcal{L}(\Lambda_{*})$ as a subset of the quotient $\mathrm{Emb}(\Lambda_{*},M)/\mathrm{Diff}(\Lambda_{*})$ of smooth embedding $\Lambda_{*}\hookrightarrow M$ by the right-action by diffeomorphisms of $\Lambda_{*}$. This quotient being endowed with the topology induced by the Whitney $C^{1}$-topology on $\mathrm{Emb}(\Lambda_{*},M)$, the $C^{1}$-topology of $\mathcal{L}(\Lambda_{*})$ is the induced topology as a subset of this quotient. Equivalently, considering $\mathcal{L}(\Lambda_{*})$ as the homogeneous space $\mathcal{G}(M,\xi)/\mathrm{Stab}(\Lambda_{*})$, its $C^{1}$-topology is the quotient topology induced by the $C^{1}$-topology of $\mathcal{G}$.
@@
Let us fix $\Lambda_{0}\in\mathcal{L}(\Lambda_{*})$. According to the Legendrian Weinstein neighborhood theorem, there exists a neighborhood $W\subset M$ of $\Lambda_{0}$ that is contactomorphic to a neighborhood of the 00-section of $J^{1}\Lambda_{0}$ endowed with the standard contact structure and that identifies $\Lambda_{0}$ with the 00-section. Every Legendrian submanifold $\Lambda$ that is $C^{1}$-close to $\Lambda_{0}$ is the $1$-jet of a map $f:\Lambda_{0}\to\mathbb{R}$. Therefore, a base of neighborhoods of the point $\Lambda_{0}$ in its isotopy class $\mathcal{L}(\Lambda_{*})$ is identifiable with the set described by the balls $B_{r}:=\{f\in C^{\infty}(\Lambda_{0},\mathbb{R})\ |\ \|f\|_{C^{1}}<r\}$, $0<r<\varepsilon$, for a small $\varepsilon>0$. In other words, the topological space $\mathcal{L}(\Lambda_{*})$ is locally modeled on the space of smooth maps on $\Lambda_{*}$ endowed with the $C^{1}$-topology. Therefore, it is a locally contractible space and it admits a genuine universal cover ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ that can be formally described as a set of equivalent classes of paths $\gamma:[0,1]\to\mathcal{L}(\Lambda_{*})$, with $\gamma(0)=\Lambda_{*}$, with the equivalence relation given by homotopy relative to endpoints. With this description of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$, the cover $\Pi:\widetilde{\mathcal{L}}\to\mathcal{L}$ is the evaluation $\gamma\mapsto\gamma(1)$.
@@
Following a similar argument applied to the contact graphs $\mathrm{gr}^{\alpha}(\psi)\subset M\times M\times\mathbb{R}$ of contactomorphisms $\psi\in\mathcal{G}(M)$, the space $\mathcal{G}(M)$ endowed with the $C^{1}$-topology is locally modeled on the space of compactly supported smooth maps $C^{\infty}_{c}(M,\mathbb{R})$ endowed with the $C^{1}$-topology (see e.g. [66, §1]). It is therefore a locally contractible space and admits a universal cover $\widetilde{\mathcal{G}}(M)$ that can be described as a space of homotopy classes. The space $\widetilde{\mathcal{G}}$ has a natural group structure making the cover $\Pi:\widetilde{\mathcal{G}}\to\mathcal{G}$ a morphism (coming from composing homotopies $(\varphi_{t})$ and $(\psi_{t})$ timewise: $(\varphi_{t}\circ\psi_{t})$). The action by conjugation of contactomorphisms of $(M,\xi)$ on $\mathcal{G}$ naturally lift to an action on $\widetilde{\mathcal{G}}$: for $g,x\in\widetilde{\mathcal{G}}$, $gxg^{-1}$ only depends on $\Pi g$ and $x$. In particular, $\Pi^{-1}\{\mathrm{id}\}$ lies in the center of $\widetilde{\mathcal{G}}$.
@@
One has a natural map $\mathrm{Cont}_{0}(M)\times\mathcal{L}(\Lambda_{*})\to\mathcal{L}(\Lambda_{*})$ given by $(\varphi,\Lambda)\mapsto\varphi(\Lambda)$. Let $I$ be an interval containing 00, if $(\varphi_{t})_{t\in I}$ is a contact isotopy with $\varphi_{0}(\Lambda_{*})=\Lambda_{*}$, its elements act naturally on ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ by defining $\varphi_{t}(\Lambda)$ for Λ~($\Lambda\in\widetilde{\mathcal{L}}(\Lambda_{*})$ and $t\in I$ as the endpoint of the path lifting $s\mapsto\varphi_{st}(\Pi\Lambda)$, $s\in[0,1]$, and starting at $\Lambda$. It induces in particular a continuous map $\widetilde{\mathcal{G}}(M)\times\widetilde{\mathcal{L}}(\Lambda_{*})\to\widetilde{\mathcal{L}}(\Lambda_{*})$.
Section 2.1 can be naturally extended to Legendrian isotopies of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$. A family $(\Lambda_{t})$ of elements of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ will be called a Legendrian isotopy if $(\Pi\Lambda_{t})$ is a Legendrian isotopy (conversely any Legendrian isotopy of $\mathcal{L}$ gives rise to a Legendrian isotopy of $\widetilde{\mathcal{L}}$ once a starting point is fixed). The contact Hamiltonian map of a Legendrian isotopy of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ will designate the contact Hamiltonian map of the projected Legendrian isotopy on $\mathcal{L}(\Lambda_{*})$ and similarly for isotopies of $\widetilde{\mathcal{G}}$. The lifting property of the universal cover implies the following extension of Lemma 2.1 (where we use the natural action of contact flows on $\widetilde{\mathcal{L}}$ discussed above).
@@
$d_{C^{1}}(\Lambda_{0},\Lambda_{1}):=\inf_{(\Lambda_{t})}\int_{0}^{1}\|H_{t}\|_{C^{1}(\Lambda_{t})}\mathrm{d}t,\quad\forall\Lambda_{0},\Lambda_{1}\in\mathcal{L}\text{ (resp. $\widetilde{\mathcal{L}}$)},$ (5)
where the infimum is taken over every isotopy $(\Lambda_{t})$ from $\Lambda_{0}$ to $\Lambda_{1}$, the Hamiltonian of which is denoted $H_{t}$, and $\|\cdot\|_{C^{1}(\Lambda)}$ is the usual $C^{1}$-norm on maps of the Riemannian submanifold $\Lambda\subset(M,g)$. The equivalence between the topology induced by $d_{C^{1}}$ and the previously described topology is a consequence of the equivalence of their base of neighborhoods which comes from the Hamilton-Jacobi equation (4).
One defines similarly a length metric on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ inducing the $C^{1}$-topology:
$d_{C^{1}}(\varphi,\psi):=\inf_{(H_{t})}\int_{0}^{1}\|H_{t}\|_{C^{1}(M,g)}\mathrm{d}t,\quad\forall\varphi,\psi\in\mathcal{G}\text{ (resp. $\widetilde{\mathcal{G}}$)},$ (6)
@@
Remark 2.4 (from $C^{1}$ to $C^{0}$).
As they induce $C^{1}$-topology on their respective space, these $C^{1}$-distances are indeed non-degenerate (i.e. only vanishing on the diagonal). However, it is not clear for their natural $C^{0}$-analogues obtained by putting $C^{0}$-norms instead of $C^{1}$-norms in their respective definition. In fact, in the case of $\mathcal{L}$, it can not be the case (see Section 2.3 below). The reason is that it is not clear which topology is associated with these metrics. Indeed, when e.g. a Legendrian $\Lambda$ is $C^{0}$-close to $\Lambda_{0}$, there is no reason for it to be a graph in any Weinstein neighborhood of $\Lambda_{0}$. Therefore, one cannot state that the $C^{0}$-topology of $\mathcal{L}$ is locally modelled on the $C^{0}$-topology of $C_{c}^{\infty}(\Lambda_{0},\mathbb{R})$. Without this local identification to a normed vector space, the induced length pseudo-metric can be degenerate or even vanish identically. Nevertheless, this is not always the case and these pseudo-distances are relevant by themselves since their non-degeneracy and geodesics are intimately linked to contact and symplectic rigidity phenomena see Section 2.3 and 4.3 below.
The following lemma will be useful in order to construct time-functions.
@@
Let $(M,\xi)$ be a contact manifold and $\Lambda_{*}\subset M$ be a closed Legendrian submanifold. The induced topological spaces $\widetilde{\mathcal{G}}(M)$ and ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ are separable.
@@
Being a subspace of the separable metrizable space of compactly supported $C^{1}$-diffeomorphisms of $M$, the space $\mathcal{G}(M)$ is a separable metrizable space. So $\mathcal{G}(M)$ admits a countable basis of open sets, which implies that $\widetilde{\mathcal{G}}(M)$ admits a countable basis of open sets by the Poincaré-Volterra theorem. Since $\widetilde{\mathcal{G}}(M)$ has a countable basis of open sets, we conclude that it is separable.
The fact that the ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ is separable is now a consequence of Lemma 2.3. ∎
@@ 2.3. Hofer type pseudo-metrics
The pseudo-distance $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ was first studied by Rosen-Zhang in [62] where it was more generally defined on non-Legendrian subsets (the equivalence of their definition with ours comes from Corollary 2.2). This pseudo-distance is known to be non-degenerate in multiple cases including when $\mathcal{L}$ is orderable [41, Theorem 5.2] (see also Corollary 4.1). However, this pseudo-distance can also be degenerate on unorderable $\mathcal{L}$ [15]. In the case it is degenerate on $\mathcal{L}$, $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ is actually identically zero, according to the generalization of the Chekanov’s dichotomy proven by Rosen-Zhang [62, Theorem 1.10].
@@
The same identities hold by considering $g,\Lambda_{0},\Lambda_{1},\varphi$ and $\psi$ in the universal cover of their respective spaces. We refer to Remark 4.6 below for a discussion on the apparent lack of invariance of these contact metrics. Topologies induced by these pseudo-distances are coarser than the $C^{1}$-topology, as they are dominated by the metric defined at (5) and (6).
@@ 2.4. Contact orderability
An isotopy $(\Lambda_{t})$ in $\mathcal{L}$ or $\widetilde{\mathcal{L}}$ is said to be non-negative (resp. positive) if its contact Hamiltonian for some, and thus for any, supporting contact form is non-negative (resp. positive). Given $\Lambda_{0},\Lambda_{1}\in\mathcal{L}$ or $\widetilde{\mathcal{L}}$, we write $\Lambda_{0}\preceq\Lambda_{1}$ (resp. $\Lambda_{0}\cll\Lambda_{1}$) if there exists a non-negative (resp. positive) isotopy $(\Lambda_{t})$ joining $\Lambda_{0}$ and $\Lambda_{1}$. One defines similarly relations $\preceq$ and on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ by considering contact Hamiltonian maps of compactly supported contact isotopies $(\varphi_{t})$, one also defines these relations on the whole space of contactomorphisms and its universal cover by considering contact isotopies.
The relation $\preceq$ is reflexive and transitive (i.e. it is a pre-order), whereas is only transitive. These relations satisfy properties of invariance with respect to the action of contactomorphisms: we have for all $x_{1},y_{1},x_{2},y_{2}\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$, resp. $\mathrm{Cont}(M,\xi)$, resp. $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$)
$\begin{cases}x_{1}\preceq y_{1}\text{ and }x_{2}\preceq y_{2}\Rightarrow x_{1}x_{2}\preceq y_{1}y_{2},\\ x_{1}\cll y_{1}\text{ and }x_{2}\preceq y_{2}\Rightarrow x_{1}x_{2}\cll y_{1}y_{2},\\ x_{1}\preceq y_{1}\text{ and }x_{2}\cll y_{2}\Rightarrow x_{1}x_{2}\cll y_{1}y_{2},\\ \end{cases}$ (11)
and the same is formally true when replacing $x_{2}$ and $y_{2}$ with elements in $\mathcal{L}$ (resp. in $\widetilde{\mathcal{L}}$).
@@
Lemma 2.7 ([30, Proposition 2.1.B], [22, Proposition 4.5]).
Let $O$ be either $\mathcal{L}$ or $-~\widetilde{\mathcal{L}}$. The space $O$ is orderable if and only if there does not exist any positive loop among isotopies of $O$. The same is true for $O$ being either $\mathcal{G}$ or $\widetilde{\mathcal{G}}$ for a closed contact manifold.
@@
Let $O$ be either $\mathcal{L}$, $-~\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$ for a given contact manifold $(M,\xi)$ and possibly a closed Legendrian submanifold $\Lambda_{*}$. Then

### v2

@@ 2.2. The $C^{1}$-topology and the universal covers
Let $\Lambda_{*}$ be a closed Legendrian submanifold of $(M,\xi)$. Let us briefly describe the $C^{1}$-topology of the space $\mathcal{L}(\Lambda_{*})$ consisting of Legendrian submanifolds $\Lambda$ that are Legendrian isotopic to $\Lambda_{*}$, i.e. there exists a Legendrian isotopy $(\Lambda_{t})$ such that $\Lambda_{0}=\Lambda_{*}$ and $\Lambda_{1}=\Lambda$, and the topology of the space $\mathcal{G}(M,\xi)$ consisting contactomorphisms $\varphi$ contact isotopic to the identity through compactly supported contactomorphisms, i.e. there exists a contact isotopy $(\varphi_{t})$ such that $\varphi_{0}=\mathrm{id}$, $\varphi_{1}=\varphi$ and $\varphi_{t}$ is compactly supported. The $C^{1}$-topology on $\mathcal{G}$ is understood as the topology induced by the strong Whitney $C^{1}$-topology on diffeomorphisms as it is described in [43, Chapter 2].
[…]
@@
Let us fix $\Lambda_{0}\in\mathcal{L}(\Lambda_{*})$. According to the Legendrian Weinstein neighborhood theorem, there exists a neighborhood $W\subset M$ of $\Lambda_{0}$ that is contactomorphic to a neighborhood of the 0-section of $J^{1}\Lambda_{0}$ endowed with the standard contact structure and that identifies $\Lambda_{0}$ with the 0-section. Every Legendrian submanifold $\Lambda$ that is $C^{1}$-close to $\Lambda_{0}$ is the 111-jet of a map $f:\Lambda_{0}\to\mathbb{R}$ which is $C^{2}$-close to the 0-map. Therefore, a base of neighborhoods of the point $\Lambda_{0}$ in its isotopy class $\mathcal{L}(\Lambda_{*})$ is identifiable with the set described by the balls $B_{r}:=\{f\in C^{\infty}(\Lambda_{0},\mathbb{R})\ |\ \|f\|_{C^{2}}<r\}$, $0<r<\varepsilon$, for a small $\varepsilon>0$. In other words, the topological space $\mathcal{L}(\Lambda_{*})$ is locally modeled on the space of smooth maps on $\Lambda_{*}$ endowed with the $C^{2}$-topology. Therefore, it is a locally contractible space and it admits a genuine universal cover ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ that can be formally described as a set of equivalent classes of paths $\gamma:[0,1]\to\mathcal{L}(\Lambda_{*})$, with $\gamma(0)=\Lambda_{*}$, with the equivalence relation given by homotopy relative to endpoints. With this description of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$, the cover $\Pi:\widetilde{\mathcal{L}}\to\mathcal{L}$ is the evaluation $\gamma\mapsto\gamma(1)$.
@@
When $M$ is closed, following a similar argument applied to the contact graphs $\mathrm{gr}^{\alpha}(\psi)\subset M\times M\times\mathbb{R}$ of contactomorphisms $\psi\in\mathcal{G}(M)$, the space $\mathcal{G}(M)$ endowed with the $C^{1}$-topology is locally modeled on the space of compactly supported smooth maps $C^{\infty}(M,\mathbb{R})$ endowed with the $C^{2}$-topology (see e.g. [67, §1]). It is therefore a locally contractible space and admits a universal cover $\widetilde{\mathcal{G}}(M)$ that can be described as a space of homotopy classes. The space $\widetilde{\mathcal{G}}$ has a natural group structure making the cover $\Pi:\widetilde{\mathcal{G}}\to\mathcal{G}$ a morphism (coming from composing homotopies $(\varphi_{t})$ and $(\psi_{t})$ timewise: $(\varphi_{t}\circ\psi_{t})$). The action by conjugation of contactomorphisms of $(M,\xi)$ on $\mathcal{G}$ naturally lift to an action on $\widetilde{\mathcal{G}}$: for $g,x\in\widetilde{\mathcal{G}}$, $gxg^{-1}$ only depends on $\Pi g$ and $x$. In particular, $\Pi^{-1}\{\mathrm{id}\}$ lies in the center of $\widetilde{\mathcal{G}}$.
Similarly, when $M$ is open we define $\widetilde{\mathcal{G}}$ (resp. $\widetilde{\mathrm{Cont}})$ as the set of equivalence classes of $C^{1}$-continuous paths $\gamma:[0,1]\to\mathcal{G}$ (resp. $\mathrm{Cont}$) starting at the identity with the equivalence relation given by homotopy relative to endpoints.
@@
One has a natural map $\mathrm{Cont}_{0}(M)\times\mathcal{L}(\Lambda_{*})\to\mathcal{L}(\Lambda_{*})$ given by $(\varphi,\Lambda)\mapsto\varphi(\Lambda)$. Let $I$ be an interval containing 0, if $(\varphi_{t})_{t\in I}$ is a contact isotopy with $\varphi_{0}(\Lambda_{*})=\Lambda_{*}$, its elements act naturally on ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ by defining $\varphi_{t}(\Lambda)$ for $\Lambda\in\widetilde{\mathcal{L}}(\Lambda_{*})$ and $t\in I$ as the endpoint of the path lifting $s\mapsto\varphi_{st}(\Pi\Lambda)$, $s\in[0,1]$, and starting at $\Lambda$. It induces in particular a continuous map $\widetilde{\mathcal{G}}(M)\times\widetilde{\mathcal{L}}(\Lambda_{*})\to\widetilde{\mathcal{L}}(\Lambda_{*})$.
2.2.1. Isotopies in $\widetilde{\mathcal{L}}$ and $\widetilde{\mathcal{G}}$ and their Hamiltonian maps
Section 2.1 can be naturally extended to Legendrian isotopies of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ and contact isotopies in $\widetilde{\mathcal{G}}$. More precisely, a family $(x_{t})$ of elements of $\widetilde{\mathcal{L}}$ (resp. $\widetilde{\mathcal{G}}$) will be called a Legendrian isotopy (resp. a contact isotopy) if $(\Pi x_{t})$ is a Legendrian isotopy (resp. a contact isotopy).
The contact Hamiltonian map of a Legendrian isotopy of ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ will designate the contact Hamiltonian map of the projected Legendrian isotopy on $\mathcal{L}(\Lambda_{*})$ and similarly for isotopies of $\widetilde{\mathcal{G}}$ (see Section 2.1).
Using the natural action of contact flows on $\widetilde{\mathcal{L}}$ discussed above we get immediatly the following extension of Lemma 2.3.
@@ 2.2.1. Isotopies in $\widetilde{\mathcal{L}}$ and $\widetilde{\mathcal{G}}$ and their Hamiltonian maps
2.2.2. Separability of $\mathcal{G},\widetilde{\mathcal{G}},\mathcal{L}$ and $\widetilde{\mathcal{L}}$
@@
$d_{C^{1}}(\Lambda_{0},\Lambda_{1}):=\inf_{(\Lambda_{t})}\int_{0}^{1}\|H_{t}\|_{C^{2}(\Lambda_{t})}\mathrm{d}t,\quad\forall\Lambda_{0},\Lambda_{1}\in\mathcal{L}\text{ (resp. $\widetilde{\mathcal{L}}$)},$ (5)
where the infimum is taken over every isotopy $(\Lambda_{t})$ from $\Lambda_{0}$ to $\Lambda_{1}$, the Hamiltonian of which is denoted $H_{t}$, and $\|\cdot\|_{C^{2}(\Lambda)}$ is the usual $C^{2}$-norm on maps of the Riemannian submanifold $\Lambda\subset(M,g)$. The equivalence between the topology induced by $d_{C^{1}}$ and the previously described topology is a consequence of the equivalence of their base of neighborhoods which comes from the Hamilton-Jacobi equation (4).
Remark 2.6.
When $M$ is closed, one defines similarly a length metric on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ inducing the $C^{1}$-topology:
$d_{C^{1}}(\varphi,\psi):=\inf_{(H_{t})}\int_{0}^{1}\|H_{t}\|_{C^{2}(M,g)}\mathrm{d}t,\quad\forall\varphi,\psi\in\mathcal{G}\text{ (resp. $\widetilde{\mathcal{G}}$)},$
@@
The following lemmas will be useful in order to construct time-functions.
@@
Let $(M,\xi)$ be a contact manifold and $\Lambda_{*}$ be a closed Legendrian submanifold. The spaces $\mathcal{L}(\Lambda_{*})$ and ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ endowed with the $C^{1}$-topology are both separable.
@@
Recall that the topological space $\mathcal{L}(\Lambda_{*})$ is seen as a subset of the base space of the projection $\mathrm{Emb}(\Lambda_{*},M)\to\mathrm{Emb}(\Lambda_{*},M)/\mathrm{Diff}(\Lambda_{*})$. The latter projection is open since $\mathrm{Diff}(\Lambda_{*})$ is a topological group acting continuously on $\mathrm{Emb}(\Lambda_{*},M)$, when both spaces are endowed with the $C^{1}$-topology. Therefore the separability of $\mathcal{L}(\Lambda_{*})$ would follow from the separability of $\mathrm{Emb}(\Lambda_{*},M)$. To prove the separability of $\mathrm{Emb}(\Lambda_{*},M)$, note that by Whitney embedding theorem one can see $\mathrm{Emb}(\Lambda_{*},M)$ as a subset of $C^{1}(\Lambda_{*},\mathbb{R}^{N})$ for $N$ large enough. Moreover $C^{1}(\Lambda_{*},\mathbb{R}^{N})$ endowed with the $C^{1}$-topology is metrizable and separable ([43, Chapter 2.1]); therefore, as a subspace of such a topological space, $\mathrm{Emb}(\Lambda_{*},M)$ is also separable, which concludes the proof of the separability of $\mathcal{L}(\Lambda_{*})$.
By (5) and what we have just said $\mathcal{L}(\Lambda_{*})$ is metrizable and separable endowed with the $C^{1}$-topology, therefore it is second-countable. Thus the Poincaré-Volterra Lemma allows to conclude that ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ is also second-countable and thus separable.∎
Lemma 2.8.
Let $(M,\xi)$ be a closed contact manifold then the spaces $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ endowed with the $C^{1}$-topology are separable.
Proof.
Since $\mathcal{G}$ can be seen as a subset of $C^{1}(M,\mathbb{R}^{N})$ for $N$ large enough by Whitney embedding theorem, one concludes as previously that $\mathcal{G}$ is separable and metrizable. Applying the Poincaré-Volterra Lemma we deduce the separability of $\widetilde{\mathcal{G}}$. ∎
@@ 2.3. Hofer type pseudo-metrics
The pseudo-distance $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ was first studied by Rosen-Zhang in [63] where it was more generally defined on non-Legendrian subsets (the equivalence of their definition with ours comes from Corollary 2.4). This pseudo-distance is known to be non-degenerate in multiple cases including when $\mathcal{L}$ is orderable [42, Theorem 5.2] (see also Corollary 4.1). However, this pseudo-distance can also be degenerate on unorderable $\mathcal{L}$ [16]. In the case it is degenerate on $\mathcal{L}$, $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ is actually identically zero, according to the generalization of the Chekanov’s dichotomy proven by Rosen-Zhang [63, Theorem 1.10].
@@
The same identities hold by considering $g,\Lambda_{0},\Lambda_{1},\varphi$ and $\psi$ in the universal cover of their respective spaces. We refer to Remark 4.6 below for a discussion on the apparent lack of invariance of these contact metrics. Topologies induced by these pseudo-distances are coarser than the $C^{1}$-topology.
@@ 2.4. Contact orderability
An isotopy $(\Lambda_{t})$ in $\mathcal{L}$ or $\widetilde{\mathcal{L}}$ is said to be non-negative (resp. positive) if its contact Hamiltonian (see Sections 2.1 and 2.2.1) for some, and thus for any, supporting contact form is non-negative (resp. positive). Given $\Lambda_{0},\Lambda_{1}\in\mathcal{L}$ or $\widetilde{\mathcal{L}}$, we write $\Lambda_{0}\preceq\Lambda_{1}$ (resp. $\Lambda_{0}\cll\Lambda_{1}$) if there exists a non-negative (resp. positive) isotopy $(\Lambda_{t})$ joining $\Lambda_{0}$ and $\Lambda_{1}$.
One defines similarly relations $\preceq$ and on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$. More precisely, a contact isotopy $(\varphi_{t})\subset\mathcal{G}$ or $\widetilde{\mathcal{G}}$, is said to be non-negative (resp. positive) if its contact Hamiltonian (as described in Sections 2.1 and 2.2.1) for some, and thus for any, supporting contact form is non-negative (resp. positive); and for any $\varphi,\psi\in\mathcal{G}$ or $\widetilde{\mathcal{G}}$, we write $\varphi\preceq\psi$ (resp. $\varphi\cll\psi$) if there exists a non-negative (resp. positive) contact isotopy $(\varphi_{t})\subset\mathcal{G}$ or $\widetilde{\mathcal{G}}$ starting at $\varphi$ and ending at $\psi$.
The relation $\preceq$ is reflexive and transitive (i.e. it is a pre-order), whereas is only transitive. As already remarked in [31] on $\widetilde{\mathcal{G}}$, these relations on $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ satisfy properties of invariance with respect to the action of contactomorphisms: we have for all $x_{1},y_{1},x_{2},y_{2}\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$)
$\begin{cases}x_{1}\preceq y_{1}\text{ and }x_{2}\preceq y_{2}\Rightarrow x_{1}x_{2}\preceq y_{1}y_{2},\\ x_{1}\cll y_{1}\text{ and }x_{2}\preceq y_{2}\Rightarrow x_{1}x_{2}\cll y_{1}y_{2},\\ x_{1}\preceq y_{1}\text{ and }x_{2}\cll y_{2}\Rightarrow x_{1}x_{2}\cll y_{1}y_{2}.\\ \end{cases}$ (10)
Proposition 2.9.
The implications stated in (10) still hold formally when $x_{1},y_{1}\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}})$ and $x_{2},y_{2}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$).
Proof.
We will only prove the first of the three implications stated in (10) since the proofs of the other implications follow the same lines.
First remark that by applying Lemma 2.3 (resp. Lemma 2.5), for any $x,y\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), we have the following equivalence
$x\preceq y\Leftrightarrow\text{ there exists }g\in\mathcal{G}\text{ (resp. }\widetilde{\mathcal{G}})\text{ such that }\mathrm{id}\preceq g\text{ and }gx=y.$ (11)
With that in hand, let us show that the following equivalence holds
$x\preceq y\Leftrightarrow\varphi x\preceq\varphi y\text{ for all }\varphi\in\mathcal{G}\text{ (resp. }\widetilde{\mathcal{G}}).$ (12)
The implication $\Leftarrow$ being obvious, we prove only the implication $\Rightarrow$. To start, suppose $x\preceq y$. Then by (11) there exists $g\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) such that $\mathrm{id}\preceq g$ and $gx=y$. Moreover by (10) $\mathrm{id}\preceq\varphi g\varphi^{-1}$ for any $\varphi\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$). Therefore applying the converse direction of (11) to $\varphi x$ we get that $\varphi x\preceq\varphi g\varphi^{-1}(\varphi x)=\varphi y$ which proves (12).
Finally take any $\varphi,\psi\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) such that $\varphi\preceq\psi$ and any $x,y\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) satisfying $x\preceq y$. So by definition $\mathrm{id}\preceq\varphi^{-1}\psi$ which implies by (11) that $x\preceq\varphi^{-1}\psi x$ and by (12) that $\varphi x\preceq\psi x$. Using (11) again, $\psi x\preceq\psi y$ and so by transitivity $\varphi x\preceq\psi y$ which concludes the proof. ∎
@@
Lemma 2.11 ([31, Proposition 2.1.B], [23, Proposition 4.5]).
Let $O$ be either $\mathcal{L}$ or $\widetilde{\mathcal{L}}$. The space $O$ is orderable if and only if there does not exist any positive loop among isotopies of $O$. The same is true for $O$ being either $\mathcal{G}$ or $\widetilde{\mathcal{G}}$ for a closed contact manifold.
@@
Let $O$ be either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$ for a given contact manifold $(M,\xi)$ and possibly a closed Legendrian submanifold $\Lambda_{*}$. Then